### v1

# An Improved Phase Coding Audio Steganography Algorithm

Abstract
As AI technology continues to advance, voice cloning is becoming increasingly easy. Recently, cases of fraud involving audio forgery using AI technology have emerged, making it particularly important to covertly embed information and verify the authenticity and integrity of audio. Digital Audio Watermarking has thus become a crucial tool in this context. This study proposes an improved Phase Coding audio steganography algorithm that dynamically segments the audio signal and embeds information into the phase components of the mid-frequency range. This approach not only enhances the algorithm’s resistance to steganalysis but also simplifies the computational process, ensuring the authenticity and integrity of audio both efficiently and securely.
Index Terms:
Audio steganography, Phase coding, Anti-steganalysis, Fast Fourier Transform, Data integrity, Bit Error Rate, Digital Audio Watermarking.
I Introduction
In recent years, cases of fraud involving AI-generated audio have become increasingly common. For instance, in 2021, some criminals used AI-generated cloned voices to impersonate the relatives or friends of victims in phone scams, leading to significant financial losses [1]. These incidents have exposed security vulnerabilities in current audio communications, particularly in areas like identity verification, digital audio watermarking and the validation of audio authenticity. Consequently, ensuring the authenticity and integrity of audio content has emerged as a critical research challenge that demands immediate attention.[2]
Embedding critical information into audio through audio steganography is a direct solution. Common audio steganography techniques include Least Significant Bit (LSB) embedding, Echo Hiding, Phase Coding, and Spread Spectrum. Among these, Phase Coding has the advantages of minimal impact on audio quality making it a favorable choice for audio steganography.[3]
Phase Coding hides information by modifying the phase components of an audio signal. This technique utilizes Fast Fourier Transform (FFT) to convert the audio signal into the frequency domain, where the secret data is embedded into the phase information. The signal is then restored to the time domain using the Inverse Fast Fourier Transform (IFFT). [4]
However, traditional Phase Coding methods have several key drawbacks:
• Low Computational Efficiency: Classic Phase Coding methods require the calculation and maintenance of phase differences, often necessitating two passes over the audio data, which increases the computational load.
• Susceptibility to Detection: The changes introduced to the phase components are often too conspicuous, making traditional Phase Coding algorithms easier to detect, thereby reducing their undetectability.[5]
• Data Integrity : The secret message is typically concentrated in certain parts of the audio, leading to uneven data distribution and challenges in ensuring overall integrity. [6]
This study aims to address the growing threat of AI-generated audio forgeries by improving the traditional Phase Coding algorithm. The goal is to enhance its undetectability, tamper resistance and computational efficiency, thereby safeguarding the authenticity and reliability of audio communications.
II Algorithm
To address the aforementioned shortcomings , this project introduces an improved Phase Coding algorithm, streamlined into five steps and designed to distribute the embedded information evenly across the entire audio signal.
1. Segmenting the Input Audio
The input signal $S$ is divided into $n$ continuous segments, $S_{i}$ ($i\in\{1,\dots,n\}$), with each segment having a length of $l$.
2. Calculating Amplitude $A_{i}$ and Phase $\phi_{i}$ for Each Segment
Using Fast Fourier Transform (FFT), the amplitude $A_{i}$ and phase $\phi_{i}$ for each segment are calculated as follows:
$A_{i}=|\text{FFT}(S_{i})|$
$\phi_{i}=\text{angle}(\text{FFT}(S_{i}))$
3. Embedding Information into the Mid-Frequency Phase
The binary data $d$ to be embedded (with a length of $m$, where $m<l$) is converted into corresponding phase values $\phi_{\text{data}}[i]$:
$\phi_{\text{data}}[i]=\begin{cases}\frac{\pi}{2},&\text{if }d_{i}=0;\\ -\frac{\pi}{2},&\text{if }d_{i}=1\end{cases}$
[…]
$\phi^{\prime}_{i}[\text{seg\_mid}-(j-1)]=\phi_{\text{data}}[j]\quad\text{for }j=1,2,\dots,\frac{m}{n}$
To maintain the odd symmetry of the Discrete Fourier Transform (DFT), the embedding operation is repeated in the corresponding mid-frequency range:
$\phi^{\prime}_{i}[\text{seg\_mid}+j]=-\phi_{\text{data}}[\frac{m}{n}-j+1]$
4. Directly Updating the Phase $\phi^{\prime}_{i}$ for Each Segment
For each segment $S_{i}$, the phase $\phi^{\prime}_{i}$ is directly updated without propagating phase differences $\Delta\phi_{i}$:
$\phi^{\prime}_{i}=\text{updated phase values}$
5. Reconstructing the Signal
The signal is reconstructed using the Inverse FFT (IFFT) on each segment $A_{i}\cdot\exp(j\phi^{\prime}_{i})$, and all segments are then concatenated to form the final audio signal:
$S^{\prime}=\text{IFFT}(A_{i}\cdot\exp(j\phi^{\prime}_{i}))$
Key Improvements of the Algorithm:
Enhanced undetectability: By embedding information into the mid-frequency range, the algorithm avoids significant phase changes in the low-frequency range, improving undetectability.
Simplified Computation: The removal of the phase difference propagation step from the traditional algorithm allows for direct phase updates within each segment, simplifying the computational process and reducing potential phase distortion.
Ensured Data Integrity: The algorithm dynamically embeds data into segments of the whole audio signal. If the audio is subjected to clipping or editing, it becomes extremely difficult to recover the original data, thus ensuring the integrity of the embedded information.
Preserved Audio Quality: the improved algorithm enhances resistance to steganalysis while maintaining the perceptual quality of the audio signal.
III Evaluation
III-A Undetectability
The improved Phase Coding algorithm embeds information into each segment of the audio, resulting in a more uniform spectral distribution that is harder to detect using spectral analysis tools.
In the phase comparison diagrams below, the blue line represents the frequency domain phase of the original audio, and the red line represents the frequency domain phase after steganography.
Figure 1: Phase Comparison of the traditional Phase Coding algorithm
In Figure 1, it is evident that the phase changes in the traditional Phase Coding method are concentrated and pronounced, making them easily detectable by tools like spectral analysis.
Figure 2: Phase Comparison of the improved Phase Coding algorithm
In contrast, the improved algorithm distributes the steganographic data more naturally, greatly enhancing its undetectability.
III-B Bit Error Rate (BER)
I tested the embedding of 8 to 1024 bits of data into a 12-second audio clip. As shown in the graph below, the blue line represents the BER for the traditional Phase Coding method, while the red line represents the BER for the improved algorithm. It can be observed that as the amount of data increases, the traditional algorithm becomes increasingly unstable, with the BER rising gradually. In contrast, the enhanced algorithm demonstrates significant improvements, with no noticeable BER until after 768 bits, and even then, the BER remains lower.
This indicates that the improved algorithm is more effective in maintaining data integrity during the embedding and extraction process, reducing the information loss or errors.
Figure 3: Bit Error Rate (BER) Comparison of Traditional and Improved Algorithms
IV Conclusion
The improved Phase Coding algorithm enhances steganography by increasing stealthiness and resistance to steganalysis, significantly lowering the bit error rate (BER). It also boosts computational efficiency and robustness against interference. With growing concerns over voice cloning and audio forgery, this algorithm provides a vital layer of security, protecting the integrity and authenticity of audio data. Future work will focus on increasing fault tolerance without adding computational complexity, particularly in safeguarding information in complex, real-world audio environments, making it crucial for applications like secure communications and identity verification.